# Performance Assessment of Lidar Odometry Frameworks: A Case Study at the Australian Botanic Garden Mount Annan

### v1

@@ 6.1 Impact of Loop Closure
The effectiveness of loop closure mechanisms is critical in mitigating cumulative drift over long trajectories. COIN-LIO’s performance indicates that its scan context algorithm built on the FAST-LIO2 framework it implements successfully detected and corrected for revisited locations.
LIO-SAM’s loop closure mechanism, while effective in smaller-scale environments, appeared less robust in this large-scale, complex setting. The divergence observed in the final quarter of the trajectory suggests that loop closures were either not detected or not effectively integrated into the optimization process. Possible reasons include insufficient feature matching due to environmental complexity or parameter settings that limited loop closure detection range.
@@ 6.2 Translation and Yaw Error Implications
The analysis of translation and yaw errors provides deeper insights into the frameworks’ performance over varying distances.
@@
Both frameworks exhibited increasing absolute translation errors with longer distances, which is expected due to the accumulation of small estimation errors over time. However, COIN-LIO maintained lower relative translation errors compared to LIO-SAM, indicating better scalability and consistency.
@@
Yaw errors increased with distance for both frameworks, but LIO-SAM showed greater angular deviation and variability, as shown in Figure 8. Accurate yaw estimation is crucial for navigation tasks that depend on precise heading information, such as path planning and obstacle avoidance. COIN-LIO’s lower yaw errors indicate a more reliable orientation estimation, enhancing its suitability for applications where angular accuracy is critical.
@@ 6.3 Robustness in Large-Scale Environments
COIN-LIO demonstrated robustness in the challenging environment of the Australian Botanic Garden. Its ability to handle asynchronous sensor data and dynamic motion contributed to maintaining trajectory accuracy. The continuous-time representation may offer advantages in environments with varying motion dynamics and sensor rates.
@@ 6.4 Limitations
- Computational Resources: The experiments were conducted on a machine with an Intel i7 7700 processor, an Nvidia RTX 3050 and 12GB of RAM. Differences in hardware could influence computational efficiency.
- Sensor Calibration: A specific metadata file to our Ouster-128 beams model was used. Any different LiDAR/IMU sensor would affect the results to some extent if the experiments had to be reproduced.
@@ 6.5 Dataset Access
The Australian Botanic Garden Mount Annan dataset is proprietary to the Australian Center for Robotics, and is not publicly available. However, the dataset can be shared upon request by contacting the authors of this paper. Interested researchers are encouraged to reach out for collaboration or data access inquiries.
@@ 6.6 Potential Improvements
For LIO-SAM, enhancing the loop closure detection range and improving feature matching algorithms could mitigate the observed drift and errors. Incorporating semantic information or employing learning-based methods for loop closure detection may also enhance robustness.
For COIN-LIO, while the performance was strong, computational efficiency could be further optimized. The optimization parameters could be adjusted to balance accuracy and processing time.
@@ 7 Conclusion
This study conducted a comprehensive evaluation of two state-of-the-art LiDAR-inertial odometry frameworks, COIN-LIO and LIO-SAM, using a challenging dataset collected in the Australian Botanic Garden. The focus was on assessing their ability to handle loop closures and mitigate cumulative drift over a long, looped trajectory.
The results demonstrated that COIN-LIO outperformed LIO-SAM in maintaining trajectory fidelity, particularly in the horizontal (x-y) plane. COIN-LIO’s trajectory estimates closely followed the ground truth, with minimal drift observed. In contrast, LIO-SAM exhibited significant drift in the x and y axes, especially in the final quarter of the trajectory, indicating challenges in its loop closure mechanism under complex environmental conditions.
Both frameworks performed comparably in the vertical (z) axis, suggesting that the IMU integration effectively supported elevation estimation. However, the divergence in horizontal accuracy highlights the importance of robust loop closure algorithms and effective sensor fusion strategies.
The analysis of relative translation and yaw errors revealed that COIN-LIO maintained lower errors and variability over varying distances compared to LIO-SAM. COIN-LIO’s consistent relative translation errors and lower yaw errors indicate better scalability and reliability for long-distance navigation tasks.
The findings suggest that COIN-LIO’s continuous-time optimization approach provide advantages in handling asynchronous sensor data and complex motion dynamics. LIO-SAM’s factor graph-based optimization, while efficient, may require enhancements to handle large-scale environments with varying feature densities.

### v2

@@ 6.1 Impact of Loop Closure
The effectiveness of loop closure mechanisms is critical in mitigating cumulative drift over long trajectories. The performance of COIN-LIO indicates that its scan context algorithm built on the FAST-LIO2 framework that it implements successfully detected and corrected for revisited locations.
LIO-SAM’s loop closure mechanism, while effective in smaller-scale environments, appeared less robust in this large-scale, complex setting. The divergence observed in the final quarter of the trajectory suggests that loop closures were not detected or not effectively integrated into the optimization process. Possible reasons include insufficient feature matching due to environmental complexity or parameter settings that limit the loop closure detection range.
@@ 6.2 Translation and Yaw Error Implications
The analysis of translation and yaw errors provides deeper insight into the performance of the frameworks over varying distances.
@@
Both frameworks exhibited an increase in absolute translation errors with longer distances, which is expected due to the accumulation of small estimation errors over time. However, COIN-LIO maintained lower relative translation errors compared to LIO-SAM, indicating better scalability and consistency.
@@
Yaw errors increased with distance for both frameworks, but LIO-SAM showed greater angular deviation and variability, as shown in Figure 8. Accurate yaw estimation is crucial for navigation tasks that depend on precise heading information, such as path planning and obstacle avoidance. The lower yaw errors of COIN-LIO indicate a more reliable orientation estimation, improving its suitability for applications where angular precision is critical.
@@ 6.3 Robustness in Large-Scale Environments
COIN-LIO demonstrated robustness in the challenging environment of the Australian Botanic Garden. Its ability to handle asynchronous sensor data and dynamic motion contributed to maintaining trajectory accuracy. Continuous-time representation may offer advantages in environments with varying motion dynamics and sensor rates.
@@ 6.4 Limitations
- Computational Resources: The experiments were carried out on a machine with an Intel i7 7700 processor, a Nvidia RTX 3050, and 12GB of RAM. Differences in hardware could influence computational efficiency.
- Sensor Calibration: A specific metadata file was used for our Ouster-128 beams model. Any different LiDAR/IMU sensor would affect the results to some extent if the experiments had to be reproduced.
@@ 6.5 Dataset Access
The Australian Botanic Garden Mount Annan dataset is proprietary to the Australian Center for Robotics and is not publicly available. However, the data set can be shared on request by contacting the authors of this paper. Interested researchers are encouraged to reach out for collaboration or data access inquiries.
@@ 6.6 Potential Improvements
For LIO-SAM, enhancing the loop closure detection range and improving the feature matching algorithms could mitigate observed drift and errors. Incorporating semantic information or employing learning-based methods for loop closure detection may also enhance robustness.
For COIN-LIO, while performance was strong, computational efficiency could be further optimized. The optimization parameters could be adjusted to balance accuracy and processing time.
@@ 7 Conclusion
This study conducted a comprehensive evaluation of two state-of-the-art LiDAR-inertial odometry frameworks, COIN-LIO and LIO-SAM, using a challenging dataset collected at the Australian Botanic Garden. The focus was on assessing their ability to handle loop closures and mitigate cumulative drift over a long, looped trajectory.
The results demonstrated that COIN-LIO outperformed LIO-SAM in maintaining trajectory fidelity, particularly in the horizontal plane (x-y). The COIN-LIO trajectory estimates closely followed the ground truth, with minimal drift observed. In contrast, LIO-SAM exhibited significant drift in the x and y axes, especially in the final quarter of the trajectory, indicating challenges in its loop closure mechanism under complex environmental conditions.
Both frameworks performed comparably on the vertical (z) axis, suggesting that the IMU integration effectively supported elevation estimation. However, the divergence in horizontal accuracy highlights the importance of robust loop closure algorithms and effective sensor fusion strategies.
Analysis of relative translation and yaw errors revealed that COIN-LIO maintained lower errors and variability over varying distances compared to LIO-SAM. The consistent relative translation errors of COIN-LIO and the lower yaw errors indicate better scalability and reliability for long-distance navigation tasks.
The findings suggest that COIN-LIO’s continuous-time optimization approach provides advantages in handling asynchronous sensor data and complex motion dynamics. LIO-SAM’s factor graph-based optimization, while efficient, may require enhancements to handle large-scale environments with varying feature densities.